The Apollo Lunar Module was used to make the transition from the spacecraft to the moon's surface and back. Consider a similar module for landing on the surface ofMars. Use conservation of mechanical energy to answer these questions.


(a) As the lander is descending, if the pilot decides to shut down the engine when the lander is at a height of 1.8 m, (this may not be a safe height to shut down theengine) and the velocity of the lander (relative to the surface of the planet) is 1.2 m/s what will be velocity of the lander at impact? Note: g on the surface of marsis about 0.4 times that on the surface of the Earth.


(b) In the case of the lunar module an impact velocity of 3.0 m/s or less was essential for a safe landing. Assuming this to be the case for the Mars lander as well,at what maximum height could the pilot shut down the engines to ensure a safe landing. Assume the velocity v0 at the time the engine is shut down is 1.2 m/s.
Answer #1
a) use X=Xo+Vot+.5at^2 to find t
then use V=Vo+at^2 to find v :D

b) same concept
answered by: MattsRiceBowl
Answer #2
A. the kinematic equation you must use is:
v^2 = u^2 +2as
now for the question we have u= 1.2 s= 1.8 and g= 9.8/4 = 2.45
now v^2 = 1.2^2 + 2*2.45*1.8 => v= root(10.26) =3.2m/s


B.Again we the same formula only we need to find s in the equation
we know that V=3m/s u =1.2m/s a= 2.45m/s^2

3^2 = 1.2^2 +2*2.45*s
=> s= (9- 1.44)/4.9 = 1.54 m
answered by: love